# A Model for Economic Freedom on Mars

### v1

@@ 2 Requirements for a Sovereign Mars
Within the constraints of this working hypothesis, Haqq-Misra [7] suggests an idealistic, although conceptually feasible, model for martian governance that would establish Mars as a sovereign planet prior to the arrival of the first humans. The purpose of such an endeavor would be to maximize the transformative potential of Mars to enable new experiments in civilization. Haqq-Misra [7] provides five specific requirements for establishing a sovereign Mars:
@@ 3 Full Reserve Banking on Mars
The model of economic freedom in this paper begins by addressing the problem of ensuring stability in the monetary system. This concept is based on the “Chicago Plan” for monetary reform that was developed by several economists during the peak of the Great Depression [11]. The Chicago Plan argued that the lack of an effective monetary authority and the widespread use of fractional reserves in banking were a major driver of the Great Depression:
@@
The major challenges and criticisms of the Chicago Plan are in its implementation on Earth today. Although Douglas et al. [11] suggested a path toward implementation, and even argued that the profession of banking could remain profitable with 100% reserves, any attempt at implementing the Chicago Plan would at the very least meet resistance from existing stakeholders that would lose their sources of wealth based on fractional reserves. Such political inertial may be difficult to overcome on Earth, although economists in more recent recessions have continued to call for reforms toward full reserve banking [e.g., 14]. But implementation would be different in the scenario of a sovereign Mars. If Mars settlement evolves in a way that is independent from such existing systems on Earth, then the development of civilization on Mars could foster novel economic experiments in full reserve banking. Space settlement itself is a long-duration ambition, so full reserve banking or a comparable sustainable monetary policy may be an inevitability for space civilization, if not on Earth.
@@ 4 Diffusion of Capital Ownership on Mars
The second component of the model of economic freedom in this paper addresses the problem of inequality in the distribution of resources. This concept is based on the proposal developed by Kelso and Adler [12] during the advent of the Cold War for the widespread dissemination of capital ownership, rather than wealth redistribution, to ensure a good life for all. Kelso and Adler [12] recognized that attempts at implementing the revolutionary ideas of Marx and Lenin by distributing wealth tended to lead to authoritarian systems that undermined individual autonomy, so they instead suggested an alternative proposal for economic reform based on the distribution of wealth-producing capital. The primary difference in these approaches is that attempts at mandating redistribution of wealth tend to enhance the centralization of government and create opportunities for authoritarianism, whereas the broad distribution of capital ownership would ensures that the production of wealth itself remains diffuse and decentralized. The concept of Kelso and Adler [12] for promoting economic freedom through the diffusion of capital ownership follows from three principles of economic justice:

### v2

@@ 2 Requirements for a Sovereign Mars
Within the constraints of this working hypothesis, Haqq-Misra [7] suggests an idealistic, although conceptually feasible, model for martian governance that would establish Mars as a sovereign planet prior to the arrival of the first humans. The purpose of such an endeavor would be to maximize the transformative potential of Mars to enable new experiments in civilization Haqq-Misra [7] provides five specific requirements for establishing a sovereign Mars:
@@ 3 Full Reserve Banking on Mars
The model of economic freedom in this paper begins by addressing the problem of ensuring stability in the monetary system. the “Chicago Plan” for monetary reform that was developed by several economists during the peak of the Great Depression [11]. The Chicago Plan argued that the lack of an effective monetary authority and the widespread use of fractional reserves in banking were a major driver of the Great Depression:
@@
The major challenges and criticisms of the Chicago Plan are in its implementation on Earth today. Although Douglas et al. [11] suggested a path toward implementation, and even argued that the profession of banking could remain profitable with 100% reserves, any attempt at implementing the Chicago Plan would at the very least meet resistance from existing stakeholders that would lose their sources of wealth based on fractional reserves. Such political may be difficult to overcome on Earth, although economists in more recent recessions have continued to call for reforms toward full reserve banking [e.g., 14]. But implementation would be different in the scenario of a sovereign Mars. If Mars settlement evolves in a way that is independent from such existing systems on Earth, then the development of civilization on Mars could foster novel economic experiments in full reserve banking. Space settlement itself is a long-duration ambition, so full reserve banking or a comparable sustainable monetary policy may be an inevitability for space civilization, if not on Earth.
@@ 4 Diffusion of Capital Ownership on Mars
The second component of the model of economic freedom in this paper addresses the problem of inequality in the distribution of resources. This concept is based on the proposal developed by Kelso and Adler [12] during the advent of the Cold War the widespread dissemination of capital ownership, wealth redistribution, to a good life for . Kelso and Adler [12] recognized that attempts at implementing the revolutionary ideas of Marx and Lenin by distributing wealth tended to lead to authoritarian systems that undermined individual autonomy, so they instead suggested an alternative proposal for economic reform based on the distribution of wealth-producing capital. The primary difference in these approaches is that attempts at mandating redistribution of wealth tend to enhance the centralization of government and create opportunities for authoritarianism, whereas the broad distribution of capital ownership would ensure that the production of wealth itself remains diffuse and decentralized.
The concept of Kelso and Adler [12] for promoting economic freedom through the diffusion of capital ownership follows from three principles of economic justice: